# Parity superselection obstructs monogamy of mutual information in free fermions

### 14 Mar 2026 v1

@@ 7 Discussion
[…]
MMI as a holographic criterion. MMI ($I_{3}\leq 0$) is a necessary condition for a state to admit a holographic dual [1, 2]. Our result shows that the same free-fermion ground state fails this test in the spin factorization ($I_{3}^{\mathrm{spin}}>0$) but passes it in the fermionic factorization ($I_{3}^{\mathrm{ferm}}<0$ for $z>z^{*}$). The outcome of the MMI test therefore depends on the operator algebra, not only on the state. This has a concrete implication: MMI violations detected in lattice numerics (DMRG, exact diagonalization) that use the tensor-product Hilbert space cannot be directly compared with holographic predictions, which implicitly assume an algebra compatible with the bulk gauge symmetry. The superselection defect identified in Proposition 1 is the precise obstruction.
@@
Open problems. Marić, Bocini, and Fagotti [13] showed that Rényi tripartite information depends on the spin structure of the underlying Thirring model; our $\Delta S_{AD}$ corresponds to the nontrivial spin-structure contribution, with the algebraic framework of [17] providing a broader context. A proof of the coherence damping inequality $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$—verified numerically for both free and interacting fermions—would imply $\Delta S_{AD}\geq 0$ for any fermionic state without Gaussian machinery.
Interacting fermions: quantitative anatomy
@@
Proposition 1 is exact for any state, not only Gaussian. For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), exact diagonalization at L=9L=9–12 confirms: (i) the coherence damping $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ holds for all tested $V$ without exception; (ii) $\Delta S_{AD}\geq 0$ for all $K$; (iii) however, $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$). For $K>K_{c}$ (free and weakly interacting): $I_{3}^{\mathrm{spin}}>0$. For $K<K_{c}$ (strong repulsion): $I_{3}^{\mathrm{spin}}<0$—both algebras satisfy MMI. The DMRG data at $L=64$ (Tables 23) confirm this picture: $I_{3}^{\mathrm{spin}}$ remains positive at $K=0.75$ for all tested $z$, consistent with $K_{c}<0.75$.
@@
This data quantifies the mechanism behind $K_{c}$: repulsion ($K<1$) suppresses parity fluctuations in $B$, reducing $\Delta I_{3}$, while simultaneously strengthening fermionic monogamy (increasing $|I_{3}^{\mathrm{ferm}}|$). At $K_{c}\approx 0.7$ these two effects cross and $I_{3}^{\mathrm{spin}}$ changes sign. The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: the ratio $R(K,z)=I_{3}(K)/I_{3}(K{=}1)$ can be decomposed as
@@
At $z=0.98$, $K=0.75$: the total deviation is 0.27-0.27, of which 0.03-0.03 is the interaction contribution and 0.24-0.24 is the factorization contribution—an $8\times$ ratio. In the fermionic basis, $R^{\mathrm{ferm}}$ stays within ${\sim}5\%$ of unity for all tested $z$ and $K$.
@@
Caveat. For $V\neq 0$, $I_{3}^{\mathrm{ferm}}$ is computed from the Gaussian formula applied to the two-point correlation matrix of a non-Gaussian state—a systematic approximation whose error is the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$. To bound this error: at $V=0$ (free fermions, Table 2), the Gaussian formula is exact, and the factorization effect $\Delta I_{3}$ is uncontaminated—this establishes the mechanism rigorously. For $V\neq 0$, exact diagonalization at $L=16$ shows that individual Gaussian entropies deviate from exact values by $\leq 1\%$ at $V=0.5$ ($K=0.86$); the non-Gaussian corrections partially cancel in the $I_{3}$ combination, so that $R^{\mathrm{ferm}}$ deviates from unity by only ${\sim}5\%$—comparable to the DMRG result at $L=64$. The near-unity of $R^{\mathrm{ferm}}$ therefore provides evidence, not proof, that the linear coefficient $c=3\ln(4/3)/\pi$ is $K$-independent. An analytical theory of how $K$ controls $K_{c}$ remains open.
@@
Operational consequences. The superselection defect has three implications. First, a single-bit measurement of $N_{B}\bmod 2$ increases I(A:D)I(A{:}D) by $\Delta S_{AD}$ (approximately 0.16 nat for $w=2$ at half filling), directly testable in cold-atom systems with parity detection [18]. Second, in JW-encoded quantum simulations, the qubit entanglement between $A$ and $D$ exceeds the fermionic entanglement by exactly $\Delta S_{AD}$, quantifying the nonlocal string overhead. Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit. Finally, cold-atom experiments measuring Rényi-2 entropy in the tensor-product basis access $I_{3}^{(2),\mathrm{spin}}$, which differs from the fermionic von Neumann $I_{3}^{(1),\mathrm{ferm}}$ analyzed in [14] in two independent respects: the Rényi index (scaling exponent $\beta=3$ vs $\beta=1$ [15]) and the operator algebra ($\Delta S_{AD}>0$, this work). Neither the sign, the zero structure, nor the Lifshitz scaling of the measured quantity can be inferred from the predictions of [14] without accounting for both effects.
@@ 8 Conclusion
The sign of $I_{3}$ is controlled by a superselection defect: the parity insertion $(-1)^{N_{B}}$ in the partial trace. Proposition 1 gives the exact identity. Theorem 1 gives the rigorous bound. Certified computations close the proof for $w=1,2$ at all fillings, extending to 2D strip geometries. For interacting fermions, the superselection defect persists and $\Delta S_{AD}\geq 0$ survives, but $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7$), delimiting the range of the obstruction. DMRG calculations at L=64L=64–128 confirm that the factorization contribution is the dominant effect at moderate $z$, accounting for ${\sim}80\%$ of the apparent $K$-dependence observed in spin-basis numerics. The outcome of the MMI test is a property of the pair (state, operator algebra), not of the state alone.

### 18 Mar 2026 v2

@@ 7 Discussion
What is proved. $I_{3}^{\mathrm{spin}}>0$ is established rigorously for: (i) all $z<z^{*}$ and all $w$ (from $g>0$ [14] and $\Delta S_{AD}\geq 0$ via Theorem 2); (ii) all $z$ at $w=1$ (analytical proof + certified computation); (iii) all $z$ at $w=2,3$ (many-body certified computation); (iv) all $z$ at $w=4,5$ (Gaussian budget from correlation matrices, Section 5.4, Table 1); (v) all $z$ at $w=6$ (many-body via $G$-matrix, Section 5.5, Fig. 2).
Conjecture 1 status. For $z<z^{*}$, $I_{3}^{\mathrm{spin}}>0$ is proved for all $w$. For $z>z^{*}$, the conjecture is proved by certified computation up to $w=6$ (Table 1, Fig. 2), and the $w=6$ data simultaneously characterize the scaling limit: $I_{3}^{\mathrm{spin}}(z)$ converges to a positive function with minimum $0.100$ at $z\approx 1.5$ (convergence $|I_{3}(w{=}6)-I_{3}(w{=}5)|<0.001$). The conjecture is therefore numerically established: the limiting function is bounded well away from zero, and finite-$w$ corrections are too small to alter the sign. A formal proof would require a rigorous bound on the convergence rate $|I_{3}(z,w)-I_{3}^{\infty}(z)|=O(1/w)$, converting the numerical evidence into a proof. Since cases (ii)–(iv) cover all $z$, the $k_{y}$-decomposition $I_{3}^{\mathrm{spin}}=\sum_{k_{y}}I_{3}^{\mathrm{spin}}(k_{F}(k_{y})\cdot w)$ [14] immediately extends the result to 2D strip geometries at $w\leq 5$.
MMI as a holographic criterion. MMI ($I_{3}\leq 0$) is a necessary condition for a state to admit a holographic dual [1, 2]. Our result shows that the same free-fermion ground state fails this test in the spin factorization ($I_{3}^{\mathrm{spin}}>0$) but passes it in the fermionic factorization ($I_{3}^{\mathrm{ferm}}<0$ for $z>z^{*}$). The outcome of the MMI test therefore depends on the operator algebra, not only on the state. This has a concrete implication: MMI violations detected in lattice numerics (DMRG, exact diagonalization) that use the tensor-product Hilbert space cannot be directly compared with holographic predictions, which implicitly assume an algebra compatible with the bulk gauge symmetry. The superselection defect identified in Proposition 1 is the precise obstruction. However, MMI is necessary but not sufficient for holography: for four adjacent blocks, the $Z_{2}$-projected entropies satisfy MMI and $I_{4}\geq 0$ but violate the holographic cyclic inequalities [19] by ${\sim}10\%$, suggesting that free-fermion entanglement does not lie in the holographic entropy cone under any of the three tested algebras (spin, fermionic, $Z_{2}$).
@@
Open problems. Marić, Bocini, and Fagotti [13] showed that Rényi tripartite information depends on the spin structure of the underlying Thirring model; our $\Delta S_{AD}$ corresponds to the nontrivial spin-structure contribution, with the algebraic framework of [17] providing a broader context. Theorem 1 settles the coherence damping inequality for ground states with real non-positive hopping; whether it extends to systems with complex hopping (e.g., in a magnetic field) or frustrated lattices remains open. In dimensions $d\geq 2$, the Jordan–Wigner string becomes path-dependent and the superselection defect acquires a geometric character absent in 1D; the algebra dependence of $I_{3}$ in higher dimensions is unexplored. All results in this paper concern gapless (metallic) free fermions; for gapped systems (e.g., the Kitaev chain), the spectral gap suppresses $\Delta S_{AD}$ and can restore $I_{3}^{\mathrm{spin}}\leq 0$ even in the spin basis—the anti-monogamy effect is a property of gapless Fermi points, not of fermionic statistics per se. For spinful fermions (e.g., the Hubbard model), $Z_{2}$ parity projection may not suffice to enforce MMI, and a finer superselection rule such as $S_{z}^{D}$ projection may be needed; this remains to be investigated.
Non-Gaussianity and full counting statistics. The Gaussian budget $\mathcal{B}$ captures only the two-point correlator contribution to $\Delta S_{AD}$; the non-Gaussian gap $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ encodes higher-order connected correlations. The JW string $(-1)^{N_{B}}=\prod_{k\in B}(1-2n_{k})$ expands into $p$-body interactions for $p=1,\ldots,w$: at $w=1$ only the one-body term contributes (hence $\varepsilon/\Delta S\approx 11\%$), while at $w=3$ terms up to $n_{1}n_{2}n_{3}$ participate ($\varepsilon/\Delta S\approx 58\%$). The Gaussian budget $\mathcal{B}$ decreases with $w$ at fixed $k_{F}$, because the spin cross-correlator $C_{ij}^{\mathrm{spin}}\propto P_{B}\sim w^{-1/2}$ (Fisher–Hartwig) decays faster than the fermionic one; for $w\gtrsim 10$, $\mathcal{B}$ turns negative (the Gaussian spin state, now nearly uncorrelated between $A$ and $D$, has higher entropy than the correlated fermionic one), making the bound $\Delta S_{AD}\geq\mathcal{B}$ trivially satisfied but uninformative. Already at $w=7$, $\mathcal{B}$ is insufficient to compensate $|g|$. Since $\varepsilon=\Delta S_{AD}$-$\mathcal{B}$ and $\mathcal{B}<0$ at large $w$, the non-Gaussian gap exceeds the total factorization defect: $\varepsilon>\Delta S_{AD}$. The Gaussian two-point correlations actually reduce $\Delta S_{AD}$ relative to the full non-Gaussian contribution; the defect survives entirely because of higher-order correlations generated by the JW string. In the scaling limit $w\to\infty$ at fixed $k_{F}$, $\Delta S_{AD}$ converges to a finite scaling function (Fig. 2) while $\mathcal{B}\to$-$\infty$, so $\varepsilon/\Delta S_{AD}\to 1$: the factorization defect becomes entirely non-Gaussian. This connects $\Delta S_{AD}$ to the full counting statistics of $N_{B}$ [22], whose cumulants $\kappa_{k}$ for $k\geq 3$ are not captured by the Gaussian approximation but contribute to the entropy at all orders. Closing Conjecture 1 for general $w$ therefore requires going beyond the Gaussian budget. The $G$-matrix formula (13) provides exactly this: it captures the full non-Gaussian $\rho_{AD}^{\mathrm{spin}}$ at polynomial cost per matrix element. The resulting scaling-limit computation (Section 5.5) bypasses the Gaussian bottleneck entirely and establishes $I_{3}^{\mathrm{spin}}>0$ at $w=6$ with a margin that exceeds the non-Gaussian gap.
8 Interacting fermions
@@
Proposition 1 is exact for any state, not only Gaussian. For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), the hopping is real and non-positive, so Theorem 1 applies and guarantees coherence damping for all $K$; exact diagonalization confirms $\Delta S_{AD}\geq 0$. Exact diagonalization at L=9L=9–12 further shows that $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$). For $K>K_{c}$ (free and weakly interacting): $I_{3}^{\mathrm{spin}}>0$. For $K<K_{c}$ (strong repulsion): $I_{3}^{\mathrm{spin}}<0$—both algebras satisfy MMI. The DMRG data at $L=64$ (Tables 23) confirm this picture: $I_{3}^{\mathrm{spin}}$ remains positive at $K=0.75$ for all tested $z$, consistent with $K_{c}<0.75$.
@@
The discrepancy $I_{3}^{\mathrm{ferm}}-g(z)$ (e.g., 0.021-0.021 at $z=1.18$) reflects $O(1/L)$ finite-size corrections, which are largest near $z^{*}$ where $g$ is small compared to $S_{AD}\sim O(1)$.
@@
This data quantifies the mechanism behind $K_{c}$: repulsion ($K<1$) suppresses parity fluctuations in $B$, reducing $\Delta I_{3}$, while simultaneously strengthening fermionic monogamy (increasing $|I_{3}^{\mathrm{ferm}}|$). At $K_{c}\approx 0.7$ these two effects cross and $I_{3}^{\mathrm{spin}}$ changes sign. The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: defining the ratio $R^{\mathrm{alg}}(K,z)=I_{3}^{\mathrm{alg}}(K)/I_{3}^{\mathrm{alg}}(K{=}1)$ separately for each algebra (alg = spin or ferm), the spin-basis deviation decomposes as
@@
At $z=0.98$, $K=0.75$: the total deviation is 0.27-0.27, of which 0.03-0.03 is the interaction contribution and 0.24-0.24 is the factorization contribution—an $8\times$ ratio. This estimate is reliable at this particular $(K,z)$ point, where the non-Gaussian gap is $\varepsilon/|I_{3}|\approx 6\%$; at larger $z$ (nearer half filling) the ratio degrades. In the fermionic basis, $R^{\mathrm{ferm}}$ stays within ${\sim}5\%$ of unity for all tested $z$ and $K$.
@@
[…]
@@
Operational consequences. The superselection defect has three implications. First, a single-bit measurement of $N_{B}\bmod 2$ increases I(A:D)I(A{:}D) by $\Delta S_{AD}$ (approximately 0.16 nat for $w=2$ at half filling), directly testable in cold-atom systems with parity detection [18]. Second, in JW-encoded quantum simulations, the qubit entanglement between $A$ and $D$ exceeds the fermionic entanglement by exactly $\Delta S_{AD}$, quantifying the nonlocal string overhead. Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7\pm 0.1$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit. Finally, cold-atom experiments measuring Rényi-2 entropy in the tensor-product basis access $I_{3}^{(2),\mathrm{spin}}$, which differs from the fermionic von Neumann $I_{3}^{(1),\mathrm{ferm}}$ analyzed in [14] in two independent respects: the Rényi index (scaling exponent $\beta=3$ vs $\beta=1$ [15]) and the operator algebra ($\Delta S_{AD}>0$, this work). Neither the sign, the zero structure, nor the Lifshitz scaling of the measured quantity can be inferred from the predictions of [14] without accounting for both effects.
The results above establish that the algebra matters: $I_{3}^{\mathrm{spin}}>0$ while $I_{3}^{\mathrm{ferm}}$ can be negative. A sharper question is whether there exists a minimal algebra restriction that enforces MMI universally. We address this next.
9 Superselection hierarchy and monogamy
[…]
$I_{3}^{\mathrm{spin}}\;>\;I_{3}^{\mathrm{ferm}}\;>\;I_{3}^{Z_{2}}\;\approx\;I_{3}^{U(1)}.$ (18)
The near-coincidence $I_{3}^{Z_{2}}\approx I_{3}^{U(1)}$ (relative difference $0.1\%$) shows that the first bit of parity information accounts for essentially all of the projection effect; the remaining $U(1)$ sectors add negligible entropy.
Three regimes emerge (Fig. 3): the spin basis gives $I_{3}>0$ for all $z$ (always “quantum”). The $Z_{2}$ and $U(1)$ projections give $I_{3}<0$ for all $z$ (always “classical”). The fermionic algebra—which twists parity-changing elements via $(-1)^{N_{B}}$ rather than projecting them to zero—is the unique algebra with a phase transition at $z^{*}$.
Figure 3: (a) $I_{3}$ under three algebras (spin, fermionic, $Z_{2}$) for $w=2$ as a function of Fermi momentum $k_{F}$. Spin (red) is always positive; $Z_{2}$ (green) always negative; fermionic (blue) changes sign at $k_{F}\approx 0.60$ (the $w=2$ zero). (b) Ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ vs $k_{F}$ for $w=1$ (solid), 2 (dashed), 3 (dotted). The dashed green line marks $2\ln 2/(3\ln\frac{4}{3})=1.606$. The ratio exceeds unity everywhere, proving $I_{3}^{Z_{2}}<0$. (c) Structure of $\rho_{AD}^{\mathrm{spin}}$ at $w=1$, $k_{F}=1$ (representative): a single off-diagonal element $\sigma=s_{2}=0.196$ (red box) is parity-changing; $Z_{2}$ projection sets it to zero. (d) Scatter of $\Delta S_{Z_{2}}$ vs $I_{3}^{\mathrm{spin}}$ for $w=1$ (circles) and $w=2$ (squares). All points lie above the diagonal (red dotted); the red-shaded region $\Delta S<I_{3}$ is empty.
Theorem 3 ($Z_{2}$ monogamy).
For free fermions with three adjacent blocks of equal width $w\leq 3$, $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}^{Z_{2}}\leq 0$ for all $z\in(0,\pi)$.
Proof.—For $w=1$ (analytical): the spin-basis $\rho_{AD}$ is $4\times 4$ with a single nonzero off-diagonal element, the parity-changing entry $\sigma=s_{2}$ (the spin cross-correlator of Sec. 6):
$\rho_{AD}^{\mathrm{spin}}=\mathrm{diag}(p_{00},p_{01},p_{01},p_{11})+\sigma\,(|01\rangle\langle 10|+|10\rangle\langle 01|),$ (19)
with $p_{00}=(1{-}n)^{2}-c_{2}^{2}$, $p_{01}=n(1{-}n)+c_{2}^{2}$, $p_{11}=n^{2}-c_{2}^{2}$, where $n=z/\pi$ and $c_{2}=\sin z\cos z/\pi$. The $Z_{2}$ projection sets $\sigma=0$. The resulting entropy increase is the concavity deficit $\Delta S_{Z_{2}}=2h_{d}(p_{01})-h_{d}(p_{01}+\sigma)-h_{d}(p_{01}-\sigma)>0$, where $h_{d}(x)=-x\ln x$. Certified computation on a grid of 20 000 points with curvature-bounded interpolation confirms $\Delta S_{Z_{2}}>I_{3}^{\mathrm{spin}}$ at all $z$ with margin $2.5\times 10^{6}$. For $w=2,3$ (certified computation): 500 grid points give margin 765 ($w=2$); 80 points give margin 13 ($w=3$). (These margins bound $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$, a different quantity from Table 1 which bounds $I_{3}^{\mathrm{spin}}$ itself.) The theorem is also verified for all $(w_{A},w_{B},w_{D})\leq 3$ (1 620 points, zero violations) and for the interacting ttVV chain at L=9L=9–12, $V\in[-1.5,2.0]$ (448 configurations, zero violations). $\square$
Remark. For $w\geq 4$ the many-body computation ($2^{3w}\times 2^{3w}$ density matrix) becomes prohibitive; however, the minimum ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ over $z$ is stable at $1.600\pm 0.002$ for $w=1,2,3$ (well above the critical value 1), and the $w$-independence of the limiting ratio (20) is proved analytically. The declining certified margins ($2.5\times 10^{6}$, $765$, $13$) reflect coarser grids at larger $w$, not a physical trend; the ratio itself shows no degradation.
Universal ratio.—At $z\to 0$, $\Delta S_{Z_{2}}\approx 2wn\ln 2$ (each of $w$ modes contributes one bit of parity entropy) while $I_{3}^{\mathrm{spin}}\approx wc\,z$ with $c=3\ln(4/3)/\pi$ [14]. The factors of $w$ cancel, giving
$\frac{\Delta S_{Z_{2}}}{I_{3}^{\mathrm{spin}}}\;\to\;\frac{2\ln 2}{3\ln(4/3)}=\frac{2}{3}\,\log_{4/3}2=1.6063\ldots$ (20)
independently of $w$, verified numerically at $w=1,2,3$ [Fig. 3(b)]. The global minimum over $z$ is 1.602 ($w=1$), 1.599 ($w=2$), 1.600 ($w=3$), all exceeding unity. The constant $4/3$ thus links the universal entanglement coefficient $c$ to a parity information bound: one Z2 bit of coherence costs 60% more entropy than the total anti-monogamy signal.
Gauge bound.—The JW string $\prod_{k\in B}\sigma_{k}^{z}=(-1)^{N_{B}}$ is a $Z_{2}$ Wilson line through $B$, with holonomy $P_{B}=\det(I-2C_{B})$ and $B$-parity sector weights $p_{\pm}=(1\pm P_{B})/2$. Projecting $\rho_{AD}^{\mathrm{spin}}$ onto $D$-parity sectors zeroes out all parity-changing elements; by the data processing inequality, the entropy increase $\Delta S_{Z_{2}}=S_{AD}^{Z_{2}}-S_{AD}^{\mathrm{spin}}$ is bounded by the entropy of the measurement outcome. Since the parity sectors of $\rho_{AD}$ are linked to those of $B$ by particle number conservation (total $N_{A}+N_{B}+N_{D}$ is fixed in each sector, so $N_{D}\bmod 2$ determines $N_{B}\bmod 2$ given $N_{A}$), this gives $\Delta S_{Z_{2}}\leq h(p_{+})$. Combining with the hierarchy $\Delta S_{AD}\leq\Delta S_{Z_{2}}$ (the fermionic twist preserves more coherence than the $Z_{2}$ projection):
$\Delta S_{AD}\;\leq\;\Delta S_{Z_{2}}\;\leq\;h\!\left(\tfrac{1+P_{B}}{2}\right)\;\leq\;\ln 2,$ (21)
where $h$ is the binary entropy. Maximum $\Delta S_{AD}/\ln 2\approx 0.27$ is reached at half filling. The bound (21) parallels the topological entanglement entropy $\gamma=\ln 2$ of the toric code [20, 21], where $Z_{2}$ superselection sectors produce the same entropic cost.
@@ 10 Conclusion
The sign of $I_{3}$ is controlled by a superselection defect: the parity insertion $(-1)^{N_{B}}$ in the partial trace. Proposition 1 gives the exact identity. Theorem 1 proves the coherence damping inequality for ground states with real non-positive hopping, implying $\Delta S_{AD}\geq 0$ for free fermions (via Theorem 2) and supporting it for interacting fermions (verified by exact diagonalization). Theorem 2 gives the Gaussian budget bound, and certified computations close the proof of $I_{3}^{\mathrm{spin}}>0$ for $w\leq 5$ at all fillings (many-body for $w\leq 3$; Gaussian budget from correlation matrices for $w=4,5$). The $G$-matrix formula (13) extends the many-body computation to $w=6$, mapping the full scaling-limit function: $I_{3}^{\mathrm{spin}}(z)$ is positive for all $z$, with a minimum of $0.100$ at $z\approx 1.5$ (Fig. 2). This numerically establishes Conjecture 1.
Theorem 3 goes further: $Z_{2}$ parity superselection on $\rho_{AD}$ enforces $I_{3}\leq 0$ at all fillings, proved for $w=1,2,3$ and verified for the interacting ttVV chain. The ratio $\Delta S_{Z_{2}}/I_{3}^{\mathrm{spin}}$ has the exact $w$-independent limiting value $2\ln 2/(3\ln\frac{4}{3})=1.606$, linking the universal coefficient $c$ of [14] to a parity information bound. The algebra defect itself is bounded by one $Z_{2}$ bit: $\Delta S_{AD}\leq h(\frac{1+P_{B}}{2})\leq\ln 2$.
For interacting fermions, $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7\pm 0.1$, the uncertainty reflecting the non-Gaussian gap in $S_{AD}$), but $I_{3}^{Z_{2}}<0$ persists at all tested $K$. DMRG calculations at L=64L=64–128 confirm that the factorization contribution dominates the apparent $K$-dependence in spin-basis numerics. The outcome of the MMI test is a property of the triple (state, partition, operator algebra), not of the state alone.